# D-CODA: Diffusion for Coordinated Dual-Arm Data Augmentation

### v1

@@ 2 Related Work
Trajectory-level augmentation methods synthesize new robot states, transitions, and/or actions. MimicGen [63], SkillMimicGen [64], and DexMimicGen [65] generate full demonstration trajectories, but rely on access to simulation or environment interaction during data generation, whereas D-CODA operates offline. Other works augment state-based inputs [66, 67] which limits applicability to vision-based learning. Zhou et al. [10] use NeRF [68] to augment visual input for corrective imitation but assume static scenes. Among the most closely related approaches is Diffusion Meets DAgger (DMD) [9], which augments single-arm eye-in-hand images with action labels using a diffusion model [69]. D-CODA builds on this foundation by demonstrating how to extend it to bimanual setups through a unified framework that synthesizes left and right wrist-camera views and employs constrained optimization to generate action labels suitable for bimanual manipulation.
@@ 3 Problem Statement and Preliminaries
We assume a bimanual robot with a left arm $l$ and right arm $r$. Throughout the following sections, mathematical notations with superscripts $l$ and $r$ denote the left and right arms, respectively. We study vision-based eye-in-hand imitation learning, which trains a policy $\pi_{\theta}$ parameterized by $\theta$ that learns from demonstration data with wrist camera images. To indicate the source arm for each wrist camera image, we use the $I^{l}$ and $I^{r}$ notation, though we may suppress the superscripts if the distinction is not necessary. To represent images at time $t$ in a demonstration, we use $I_{t}^{l}$ and $I_{t}^{r}$. All images are in $\mathbb{R}^{H\times W\times 3}$ with matching height $H$ and width $W$ values. These form the policy input, which produces actions ${\mathbf{a}_{t}=\pi_{\theta}((I_{t}^{l},I_{t}^{r}))}$. Here, ${\mathbf{a}_{t}=(\mathbf{a}_{t}^{l},\mathbf{a}_{t}^{r})}$, where $\mathbf{a}_{t}^{l}$ and $\mathbf{a}_{t}^{r}$ are target joint positions for the respective arms. To train $\pi_{\theta}$, imitation learning uses a dataset of expert demonstrations ${\mathcal{D}=\{\tau_{1},\ldots,\tau_{M}\}}$. Each $\tau_{i}$ is a sequence of wrist-camera images observations and actions: ${\tau_{i}=(I_{1}^{l},I_{1}^{r},\mathbf{a}_{1}^{l},\mathbf{a}_{1}^{r},\ldots,\mathbf{a}_{T-1}^{l},\mathbf{a}_{T-1}^{r},I_{T}^{l},I_{T}^{r})}$ for a demonstration with $T$ time steps.
@@ 4.1 Diffusion Model for Novel-View Synthesis
and where $z_{b,0}^{l}=V(I_{b}^{l})$ and $z_{b,0}^{r}=V(I_{b}^{r})$. The diffusion model architecture [69] is based on U-Net [72], which consists of convolution, cross-attention, and self-attention layers. To condition the model on pose transformations, we inject $\Delta p^{l},(\Delta p^{l})^{-1},\Delta p^{r},(\Delta p^{r})^{-1}$ into the cross attention layers. This improves the feature representations by incorporating relative camera pose information between the source and target views [69]. During training, we randomly sample images $\{I_{a}^{l},I_{b}^{l},I_{a}^{r},I_{b}^{r}\}$ from a robot trajectory to construct the input $(I_{a}^{l},I_{b}^{l},\Delta p^{l},I_{a}^{r},I_{b}^{r},\Delta p^{r})$ for the model, and we compute $\Delta p={}_{a}T_{b}$ by taking the matrix product of the inverse of camera pose $a$ and camera pose $b$. Given a dataset of expert demonstrations $\mathcal{D}$, we train the diffusion model on $\mathcal{D}$ for a fixed number of iterations. We then use the trained model and sampled camera perturbations (subsection 4.2) to synthesize novel wrist camera views based on the same dataset.
@@ 4.2 Camera Perturbation Sampling
If gripper-to-object contact is absent, we uniformly sample a random direction for each arm within a predefined range of magnitudes [$m_{lb}$, $m_{ub}$] and rotations [$r_{lb}$, $r_{ub}$] to generate camera perturbations. However, when contact is detected, we formulate camera perturbation sampling as a constrained optimization problem. Our key insight is to apply identical perturbations to both arms during contact events, ensuring coordinated behavior. For this, we employ Dual Annealing [74], a global optimization algorithm capable of handling constraints, with early stopping. The decision variables are the translation coordinates $c_{\rm trans}$, representing the transformation applied to the camera perturbations (normalized to [-1, 1]). The cost function penalizes perturbations that are too small, and end-effector poses that are either too close to the table or too close to the other end-effector. Additionally, we use an inverse kinematics solver based on the Levenberg-Marquardt (LM) method to check the feasibility of the perturbed end-effector poses. We define the overall optimization problem as:
$\underset{c_{\text{trans}}}{\rm minimize}$ $\text{Cost}(c_{\text{trans}})\quad\text{subject to}\quad\begin{cases}c_{\text{trans}}\in[-1,1]^{3}\quad\mbox{and}\quad c_{\text{trans}}\geq m_{\text{lb}}\\ \text{ProximityToTable}(c_{\text{trans}})\geq d_{\text{table}}\\ \text{ProximityToOtherEEF}(c_{\text{trans}})\geq d_{\text{eff}}\\ \text{IKSolver}(c_{\text{trans}})=\text{valid}\end{cases}$
@@ 4.3 Action Labeling and Dataset Construction
Given the dataset $\mathcal{D}$ and its corresponding sampled camera perturbations, we use the trained diffusion model to synthesize novel wrist camera views for both arms. To generate perturbed end-effector poses, we perform matrix multiplications involving the camera perturbation transformation $T$, the original camera pose $C$, and the end-effector pose $E$: $C\cdot T\cdot(C)^{-1}\cdot E$. Since our eye-in-hand imitation learning algorithm operates in joint space, we use the LM inverse kinematics solver to compute the perturbed target joint positions $\tilde{\mathbf{a}}_{t}$ ($\tilde{\mathbf{a}}_{t}^{l}$ and $\tilde{\mathbf{a}}_{t}^{r}$). If the resulting configuration is invalid, we discard the augmentation for that state and retain the original state information. Otherwise, we replace the original state with the augmented state every $k$ timesteps. This will result in an augmented dataset of novel views $\tilde{\mathcal{D}}$, which we combine with the original dataset, and $\pi_{\theta}$ is trained on $\mathcal{D}=\mathcal{D}\cup\tilde{\mathcal{D}}$. See Figure 3 for a visualization of the original and augmented camera positions in a combined dataset, and Figure 4 for examples of the synthesized images.
@@ 5 Experiments and Results
[…]
Figure 4: Examples of the original and synthesized wrist-camera images from both arms using D-CODA on Coordinated Lift Ball and Coordinated Lift Tray tasks in simulation. The first black column of images are the original states where the following column of red images are the augmented (perturbed original) states. All original and augmented state pairs are at the same timestep and each task is from the same episode.
@@ 5.4 Real-World Results
Figure 5: Example successful rollouts (top: Lift Drawer; bottom: Push Block) of D-CODA on a real-world bimanual setup with UR5s. See Section 5.4 for quantitative results.
@@
The real-world experiment results are in Table 2, with example rollouts in Figure 5. Fine-tuned VISTA uses all three cameras, following its best-performing variant, whereas the other methods use only the wrist cameras, except in Push Block, where we found that all methods benefit from third-person views. D-CODA outperforms baselines on all three tasks based on evaluations over 20 trials. We observe that in Lift Ball, the robot arms freeze less frequently with D-CODA compared to baselines when the arms are in contact and lifting the ball. In Push Block, they get stuck less often when the block is positioned farther from the grippers, compared to the ACT baseline. In Lift Drawer, our method reaches the sides of the drawer more frequently than the baselines. Fine-tuned VISTA performs very poorly on this task, similar to its performance in Coordinated Lift Tray Easy. In 9 out of 20 trials, VISTA successfully reached the sides of the drawer, but the policy failed to close the grippers. These results suggest that VISTA struggles with tasks requiring precise manipulation, as shown in both simulation and real-world experiments. We suspect that this limitation arises from the use of augmentations for third-person views, which may adversely affect policy learning when wrist-camera views are more critical for task success. In other words, the policy may prioritize learning invariant features from third-person views rather than focusing on task-relevant features in the wrist-camera views.
@@ 7 Limitations
Appendix A Task Details
@@
Figure 7: Real-world bimanual manipulation tasks.
@@ Appendix B Ablations
Appendix C Real-World $\pi_{0}$-FAST Experiment
Method Lift Ball Lift Drawer Push Block
$\pi_{0}$-FAST (w/o augment.) 2 / 20 1 / 20 20 / 20
D-CODA (ours) 12 / 20 1 / 20 20 / 20
Table 4: Real-world experiment results comparing $\pi_{0}$-FAST without data augmentation and $\pi_{0}$-FAST with D-CODA, with 20 trials per method and task combination.
We fine-tune a vision-language-action model, $\pi_{0}$-FAST [81], on the original and augmented data described in Section 5. $\pi_{0}$-FAST (w/o augment.) is trained only on the original data, whereas D-CODA (ours) is trained on both the original and augmented data. Fine-tuning is performed using Low-Rank Adaptation (LoRA) with the Gemma-2B-LoRA variant for 150,000 training steps, provided in [81]. In Lift Ball, the baseline frequently misses the ball by moving the arms over it, or squeezes the ball so tightly that it triggers a force limit error on the robot. In contrast, D-CODA more reliably completes the task by positioning the arms beneath the ball to lift it, a strategy that is not observed in the baseline executions. However, most failures of D-CODA are due to large action values generated from the policy, causing the arms to deviate from the intended trajectory. We suspect that the large actions may result from the discontinuous nature of action tokens, as the augmented states are out-of-distribution, perturbed original states, which could inadvertently cause the policy to learn to output actions that suddenly deviate from the trajectory. This issue does not appear in ACT and might be mitigated by adopting a smoother action token representation. In Lift Drawer, both D-CODA and the baseline achieve the same success rate. However, in 12 out of 20 trials, D-CODA has at least one gripper reach the side of the drawer (an intermediate subgoal necessary to complete the task), compared to only 9 out of 20 trials for the baseline. In Push Block, both D-CODA and the baseline succeed in 20 out of 20 trials. Overall, both $\pi_{0}$-FAST and ACT demonstrate improved performance with our data augmentation; however, with limited training data, ACT appears to exhibit greater reliability.
@@ Appendix D Generalization Experiment
Table 5: Ablation experiment results in simulation.
@@ Appendix E Additional Implementation Details
For training the diffusion model, we use the same VQ-GAN pre-trained checkpoint at 2000 epochs as DMD [9]. To randomly sample images ${I_{a}^{l},I_{b}^{l},I_{a}^{r},I_{b}^{r}}$ from a robot trajectory to construct the input $(I_{a}^{l},I_{b}^{l},\Delta p^{l},I_{a}^{r},I_{b}^{r},\Delta p^{r})$, we sample from a range of $\{5,\ldots,15\}$ for all simulation tasks, except for Dual Push Buttons, where we use $\{10,\ldots,30\}$. For real-world experiments, we use a range of $\{1,\ldots,3\}$ for all tasks. For example, if ${I_{a}^{l},I_{a}^{r}}$ are from timestep $t$, then ${I_{b}^{l},I_{b}^{r}}$ are sampled from a future timestep between $t+1$ and $t+3$ in real-world tasks. We use two NVIDIA 4060 Ti GPUs for both training the diffusion model and performing image synthesis.
For camera perturbation sampling, the translation magnitudes [$m_{lb}$, $m_{ub}$] are set to 0.01 and 0.02 meters, respectively, for contactless and contact-rich states. For contactless states, the rotation bounds [$r_{lb}$, $r_{ub}$] are set to –28.7 and 28.7 degrees, respectively. For $k$ (i.e., the interval at which original states are replaced), we set $k=6$ for Coordinated Lift Ball, Coordinated Lift Tray Easy, Dual Push Buttons, and all real-world tasks, and $k=9$ for Coordinated Push Box Easy and Bimanual Straighten Rope. Figure 9 shows that our method is largely insensitive to the choice of $k$, which motivates our choice of 6 as the default value for most tasks.

### v2

@@ 2 Related Work
Trajectory-level augmentation methods [63, 64, 65, 66] synthesize new robot states, transitions, and/or actions. MimicGen [64], SkillMimicGen [65], and DexMimicGen [66] generate full demonstration trajectories, but rely on access to simulation or environment interaction during data generation, whereas D-CODA operates offline. Other works augment state-based inputs [67, 68] which limits applicability to vision-based learning. Zhou et al. [10] use NeRF [69] to augment visual input for corrective imitation but assume static scenes. Among the most closely related approaches is Diffusion Meets DAgger (DMD) [9], which augments single-arm eye-in-hand images with action labels using a diffusion model [70]. D-CODA builds on this foundation by demonstrating how to extend it to bimanual setups through a unified framework that synthesizes left and right wrist-camera views and employs constrained optimization to generate action labels suitable for bimanual manipulation.
@@ 3 Problem Statement and Preliminaries
We assume a bimanual robot with a left arm $l$ and right arm $r$. Throughout the following sections, mathematical notations with superscripts $l$ and $r$ denote the left and right arms, respectively. We study vision-based eye-in-hand imitation learning, which trains a policy $\pi_{\theta}$ parameterized by $\theta$ that learns from demonstration data with wrist camera images. To indicate the source arm for each wrist camera image, we use the $I^{l}$ and $I^{r}$ notation, though we may suppress the superscripts if the distinction is not necessary. To represent images at time $t$ in a demonstration, we use $I_{t}^{l}$ and $I_{t}^{r}$. All images are in $\mathbb{R}^{H\times W\times 3}$ with matching height $H$ and width $W$ values. These form the policy input, which produces actions $\mathbf{a}_{t}=\pi_{\theta}((I_{t}^{l},I_{t}^{r}))$}. Here, $\mathbf{a}_{t}=(\mathbf{a}_{t}^{l},\mathbf{a}_{t}^{r})$}, where $\mathbf{a}_{t}^{l}$ and $\mathbf{a}_{t}^{r}$ are target joint positions for the respective arms. To train $\pi_{\theta}$, imitation learning uses a dataset of expert demonstrations $\mathcal{D}=\{\tau_{1},\ldots,\tau_{M}\}$}. Each $\tau_{i}$ is a sequence of wrist-camera images observations and actions: $\tau_{i}=(I_{1}^{l},I_{1}^{r},\mathbf{a}_{1}^{l},\mathbf{a}_{1}^{r},\ldots,I_{T}^{l},I_{T}^{r},\mathbf{a}_{T}^{l},\mathbf{a}_{T}^{r})$} for a demonstration with $T$ time steps.
@@ 4.1 Diffusion Model for Novel-View Synthesis
and where $z_{b,0}^{l}=V(I_{b}^{l})$ and $z_{b,0}^{r}=V(I_{b}^{r})$. The diffusion model architecture [70] is based on U-Net [73], which consists of convolution, cross-attention, and self-attention layers. To condition the model on pose transformations, we inject $\Delta p^{l},(\Delta p^{l})^{-1},\Delta p^{r},(\Delta p^{r})^{-1}$ into the cross-attention layers. This improves the feature representations by incorporating relative camera pose information between the source and target views [70]. During training, we randomly sample images $\{I_{a}^{l},I_{b}^{l},I_{a}^{r},I_{b}^{r}\}$ from a robot trajectory to construct the input $(I_{a}^{l},I_{b}^{l},\Delta p^{l},I_{a}^{r},I_{b}^{r},\Delta p^{r})$ for the model, and we compute $\Delta p={}_{a}T_{b}$ by taking the matrix product of the inverse of camera pose $a$ and camera pose $b$. Given a dataset of expert demonstrations $\mathcal{D}$, we train the diffusion model on $\mathcal{D}$ for a fixed number of iterations. We then use the trained model and sampled camera perturbations (subsection 4.2) to synthesize novel wrist camera views based on the same dataset.
@@ 4.2 Camera Perturbation Sampling
If gripper-to-object contact is absent, we uniformly sample a random direction for each arm within a predefined range of magnitudes $m_{lb}$, $m_{ub}$] and rotations $r_{lb}$, $r_{ub}$] to generate camera perturbations. However, when contact is detected, we formulate camera perturbation sampling as a constrained optimization problem. Our key insight is to apply identical perturbations to both arms during contact events, ensuring coordinated behavior. For this, we employ Dual Annealing [75], a global optimization algorithm capable of handling constraints, with early stopping. The decision variables are the translation coordinates $c_{\rm trans}$, representing the transformation applied to the camera perturbations (normalized to [-1, 1]). The cost function penalizes perturbations that are too small, and end-effector poses that are either too close to the table or too close to the other end-effector. Additionally, we use an inverse kinematics solver based on the Levenberg-Marquardt (LM) method to check the feasibility of the perturbed end-effector poses. We define the overall optimization problem as:
[…]
@@ 4.3 Action Labeling and Dataset Construction
Given the dataset $\mathcal{D}$ and its corresponding sampled camera perturbations, we use the trained diffusion model to synthesize novel wrist camera views for both arms. To generate perturbed end-effector poses, we perform matrix multiplications involving the camera perturbation transformation $T$, the original camera pose $C$, and the end-effector pose $E$: $C\cdot T\cdot(C)^{-1}\cdot E$. Since our eye-in-hand imitation learning algorithm operates in joint space, we use the LM inverse kinematics solver to compute the perturbed target joint positions $\tilde{\mathbf{a}}_{t}$ ($\tilde{\mathbf{a}}_{t}^{l}$ and $\tilde{\mathbf{a}}_{t}^{r}$). If the resulting configuration is invalid, we discard the augmentation for that state and retain the original state information. Otherwise, we replace the original state with the augmented (out-of-distribution) state every $k$ timesteps, which helps mitigate the issue of compounding errors in behavior cloning policies. The non-augmented action labels and corresponding states remain in-distribution, which guides the behavior cloning policies to complete the tasks. This will result in an augmented dataset of novel views $\tilde{\mathcal{D}}$, and $\pi_{\theta}$ is trained on $\mathcal{D}=\mathcal{D}\cup\tilde{\mathcal{D}}$. See Figure 3 for a visualization of the original and augmented camera positions in a combined dataset, and Figure 4 for examples of the synthesized images.
@@ 5 Experiments and Results
  • Coordinated Lift Ball: a ball is randomly spawned in a workspace of $0.65\times 0.91$  $\text{\,}\mathrm{m}$, same as PerAct2. A success is when the ball is lifted to a height above $0.95\text{\,}\mathrm{m}$.
  • Coordinated Lift Tray Easy: a tray with an item placed on top is randomly spawned in a workspace of $0.46\times 0.64$  $\text{\,}\mathrm{m}$, 70% of PerAct2’s original workspace, and the tray does not rotate. A success is when the tray and the item are lifted to a height above $1.2\text{\,}\mathrm{m}$.
  • Coordinated Push Box Easy: a large box and a target area are randomly spawned in a workspace of $0.59\times 0.82$  $\text{\,}\mathrm{m}$, 90% of PerAct2’s original workspace, and the box may be randomly rotated by up to 4 degrees. A success is when the box reaches the target area.
  • Dual Push Buttons: three buttons with different colors are randomly spawned in a workspace of $0.65\times 0.91$  $\text{\,}\mathrm{m}$, as in PerAct2. A success is when two specified buttons are pressed simultaneously.
  • Bimanual Straighten Rope: a long rope and target areas are randomly spawned in a workspace of $0.65\times 0.91$  $\text{\,}\mathrm{m}$, same as PerAct2. A success is when both ends of a rope are in their target areas.
Real-World. We use three coordinated tasks: Lift Ball, Lift Drawer, and Push Block, as coordinated dual-arm data augmentation is our primary focus. We use two CB2 UR5 6-DOF robot arms in a bimanual setup in a $0.97\times 0.79$  $\text{\,}\mathrm{m}$ workspace, with a front camera and two wrist-mounted cameras. Each arm has a Robotiq 2F-85 parallel-jaw gripper and an Intel RealSense D415 RGB-D wrist camera. An experienced roboticist uses GELLO [78] to teleoperate the robots and collects $\sim$32 demonstrations per task. For evaluation, we perform 20 rollouts per task. In Lift Ball, a $0.35\text{\,}\mathrm{m}$ diameter ball is placed in randomized positions within a $0.64\times 0.20$  $\text{\,}\mathrm{m}$ region. A success is when the ball is lifted to a height above $0.25\text{\,}\mathrm{m}$. In Lift Drawer, a $0.29\times 0.29\times 0.15$  $\text{\,}\mathrm{m}$ square drawer is placed randomly within a $0.48\times 0.38$  $\text{\,}\mathrm{m}$ region and randomly rotated up to 25°. A success is when the drawer is lifted to a height above $0.22\text{\,}\mathrm{m}$. In Push Block, a $0.07\times 0.35\times 0.12$  $\text{\,}\mathrm{m}$ foam block is randomly placed within a $0.97\times 0.43$  $\text{\,}\mathrm{m}$ region and rotated up to 13°. A success is when the block is pushed past the front of the workspace.
Figure 4: Examples of the original and synthesized wrist-camera images using D-CODA on Coordinated Lift Ball and Coordinated Lift Tray tasks in simulation. The first black column of images are the original states where the following column of red images are the augmented (perturbed original) states. All original and augmented state pairs are at the same timestep and each task is from the same episode. See  Appendix H for more examples.
@@ 5.4 Real-World Results
Real-world results are shown in Table 2 and example rollouts in Figure 6. The top three rows use ACT as the downstream manipulation policy, while the bottom two use $\pi_{0}$-FAST [80], a vision-language-action model. $\pi_{0}$-FAST is fine-tuned using Low-Rank Adaptation (LoRA) with the Gemma-2B-LoRA variant for 150,000 training steps, provided in [80]. Fine-tuned VISTA uses all three cameras, following its best-performing variant, whereas the other methods use only the wrist cameras, except in Push Block, where we found that all methods benefit from third-person views. D-CODA outperforms baselines on all three tasks based on evaluations over 20 trials.
We observe that in Lift Ball, when using ACT, the robot arms freeze less frequently with D-CODA compared to baselines when the arms are in contact and lifting the ball. When using $\pi_{0}$-FAST, the baseline frequently misses the ball by moving the arms over it, or squeezes the ball so tightly that it triggers a force limit error on the robot. In contrast, D-CODA more reliably completes the task by positioning the arms beneath the ball to lift it, a strategy not seen in the baseline. However, most failures of D-CODA are due to large action values generated from the policy, causing the arms to deviate from the intended trajectory. We suspect that the large actions may result from the discontinuous nature of action tokens, as the augmented states are out-of-distribution, perturbed original states, which could inadvertently cause the policy to learn to output actions that suddenly deviate from the trajectory. This issue does not appear in ACT and might be mitigated by adopting a smoother action token representation. In Push Block, the robot arms using D-CODA get stuck less often when the block is positioned farther from the grippers, compared to the ACT baseline. In Lift Drawer, our method reaches the sides of the drawer more frequently than the baselines, an intermediate subgoal necessary to complete the task. Fine-tuned VISTA performs very poorly on this task, similar to its performance in Coordinated Lift Tray Easy. In 9 out of 20 trials, VISTA successfully reached the sides of the drawer, but the policy failed to close the grippers. These results suggest that VISTA struggles with tasks requiring precise manipulation, as shown in both simulation and real-world experiments. We suspect that this limitation arises from the use of augmentations for third-person views, which may adversely affect policy learning when wrist-camera views are more critical for task success. In other words, the policy may prioritize learning invariant features from third-person views rather than focusing on task-relevant features in the wrist-camera views. Overall, both $\pi_{0}$-FAST and ACT demonstrate improved performance with our data augmentation; however, with limited training data, ACT appears to exhibit greater reliability.
@@ 7 Limitations
8 Acknowledgments
We thank our colleagues from the Robotic Embedded Systems Laboratory (RESL) and the Sensing, Learning, and Understanding for Robotic Manipulation (SLURM) lab at the University of Southern California for their fruitful discussions and helpful writing feedback.
Appendix A Paper Changelog
Version 1 on arXiv was the initial public release of the paper. Version 2 included minor writing and BibTeX edits and served as the camera-ready version for CoRL 2025.
Appendix B Task Details
@@
Figure 6: Top: Real-world bimanual manipulation tasks. Bottom: Example successful rollouts (Lift Drawer on top row; Push Block on bottom row) of D-CODA on a real-world bimanual setup with UR5s. See Section 5.4 for quantitative results.
@@ Appendix D Generalization Experiment
Table 4: Generalization experiment results in simulation.
@@ Appendix E Additional Implementation Details
For training the diffusion model, we use the same VQ-GAN pre-trained checkpoint at 2000 epochs with frozen codebooks as DMD [9]. To randomly sample images $I_{a}^{l},I_{b}^{l},I_{a}^{r},I_{b}^{r}$} from a robot trajectory to construct the input $(I_{a}^{l},I_{b}^{l},\Delta p^{l},I_{a}^{r},I_{b}^{r},\Delta p^{r})$, we sample from a range of $\{5,\ldots,15\}$ for all simulation tasks, except for Dual Push Buttons, where we use $\{10,\ldots,30\}$. For real-world experiments, we use a range of $\{1,\ldots,3\}$ for all tasks. For example, if $I_{a}^{l},I_{a}^{r}$} are from timestep $t$, then $I_{b}^{l},I_{b}^{r}$} are sampled from a future timestep between $t+1$ and $t+3$ in real-world tasks. We use two NVIDIA 4060 Ti GPUs for both training the diffusion model and performing image synthesis.
@@
For camera perturbation sampling, the translation magnitudes $m_{lb}$, $m_{ub}$] are set to 0.01 and 0.02 meters, respectively, for contactless and contact-rich states. For contactless states, the rotation bounds $r_{lb}$, $r_{ub}$] are set to –28.7 and 28.7 degrees, respectively. For $k$ (i.e., the interval at which original states are replaced), we set $k=6$ for Coordinated Lift Ball, Coordinated Lift Tray Easy, Dual Push Buttons, and all real-world tasks, and $k=9$ for Coordinated Push Box Easy and Bimanual Straighten Rope. Figure 8 shows that our method is largely insensitive to the choice of $k$, which motivates our choice of 6 as the default value for most tasks.